The student mixes 0.0140 g of CaO solids with 0.02
Answer #1

CaO + Na2CO3 ------ CaCO3 + Na2O

Molar mass of CaO = 40 + 16 = 56 gm/mol

Number of moles of CaO = 0.0140/56 = 2.5 * 10^(-4)

Molar mass of Na2CO3 = 23 * 2 + 12 + 48 = 106 gm/mol

Number of moles of CaO = 0.0265/106 = 2.5 * 10^(-4)

Q = [Ca+2][CO3(2-)] = (2.5 * 10^(-4))^2 = 6.25 * 10^(-8) > Ksp

Ksp of CaCO3 will be 3.3 * 10^(-9)

Hence the precipiate will be formed

b)

number of moles of CaCO3 formed = 2.5 * 10^(-4) moles

Molar mass of CaCO3 = 40 + 12 + 3 * 16 = 100 gm/mol

Mass of CaCO3 formed = 2.5 * 10^(-4) mol * 100 gm/mol = 2.5 * 10^(-2) gms